### v1

# POMATO: Marrying Pointmap Matching with Temporal Motions for Dynamic 3D Reconstruction

Abstract
3D reconstruction in dynamic scenes primarily relies on the combination of geometry estimation and matching modules where the latter task is pivotal for distinguishing dynamic regions which can help to mitigate the interference introduced by camera and object motion. Furthermore, the matching module explicitly models object motion, enabling the tracking of specific targets and advancing motion understanding in complex scenarios. Recently, the proposed representation of pointmap in DUSt3R suggests a potential solution to unify both geometry estimation and matching in 3D space, but it still struggles with ambiguous matching in dynamic regions, which may hamper further improvement. In this work, we present POMATO, a unified framework for dynamic 3D reconstruction by marrying POintmap MAtching with Temporal mOtion. Specifically, our method first learns an explicit matching relationship by mapping RGB pixels from both dynamic and static regions across different views to 3D pointmaps within a unified coordinate system. Furthermore, we introduce a temporal motion module for dynamic motions that ensures scale consistency across different frames and enhances performance in tasks requiring both precise geometry and reliable matching, most notably 3D point tracking. We show the effectiveness of the proposed pointmap matching and temporal fusion paradigm by demonstrating the remarkable performance across multiple downstream tasks, including video depth estimation, 3D point tracking, and pose estimation. Code and models are publicly available at https://github.com/wyddmw/POMATO.
Figure 1: 3D reconstruction from an arbitrary dynamic video with POMATO. Without relying on external modules, POMATO can directly perform 3D reconstruction along with temporal 3D point tracking and dynamic mask estimation.
{NoHyper}footnotetext: * Equal contribution. 🖂 Corresponding author.
1 Introduction
Figure 2: Ambiguity in 3D point matching in dynamic scenes with DUSt3R. Given representative corresponding pixels of background (orange) and moving foreground (red) in two different views, DUSt3R outputs a pair of 3D points within the same coordinate system. In static regions, identical pixels share the same 3D coordinates which provide an accurate matching relationship in 3D space, but in moving regions, the 3D coordinates are inconsistent for corresponding pixels across views, leading to ambiguous 3D matching relationships.
Image-based 3D reconstruction is a fundamental task in computer vision with a wide range of applications including SLAM [37], robotics [17, 47], autonomous driving [51], and novel view synthesis [5]. While substantial progress has been achieved in static 3D reconstruction [14, 49, 21, 42, 24], dynamic scenes remain significantly more challenging. The presence of moving objects introduces random motion and deformation, which can interfere with the learning of local structure and camera motion, thus complicating accurate geometry estimation. These scenarios require joint modeling of both scene geometry and object motion. Moreover, downstream tasks such as 3D point tracking demand precise geometry estimation and robust matching across views. To effectively distinguish dynamic regions, it is essential to establish reliable correspondences between frames. Some pioneering works have attempted to address dynamic motion by incorporating auxiliary matching modules, such as optical flow [40, 50] or 2D tracking [45]. However, these approaches may suffer from domain gaps and accumulated errors between modules, limiting their effectiveness. A unified framework that seamlessly integrates geometry estimation and matching for dynamic 3D reconstruction remains a critical and underexplored challenge.
Recently, DUSt3R [42] proposes a promising solution to address this challenge. It introduces the concept of a pointmap that assigns each pixel in an image to a corresponding 3D coordinate. The network utilizes a standard transformer-based encoder-decoder architecture and receives a pair of images as input. The system incorporates two parallel decoders with a regression head to predict pointmaps for each view within the same coordinate system as the initial view. However, the representation of the pointmap still faces limitations when handling ambiguous correspondence in dynamic scenes, as illustrated in Fig. 2. In static regions, the pointmap of the second image accurately reflects the rigid camera transformation from the first view, where identical RGB pixels correspond to the same 3D coordinates (orange point). Conversely, in dynamic regions, pixel displacements or deformations due to object motion introduce inconsistencies in the predicted pointmaps across views (red point), resulting in inaccurate 3D correspondences. This inconsistency fails to depict the accurate matching relationships in 3D space.
To tackle this problem, we present POMATO, a unified network for dynamic 3D reconstruction by marrying POintmap MAtching with Temporal mOtion. We introduce an auxiliary matching head that leverages decoder tokens from the second image while preserving matching features through iterative cross-attention across views. This module is tasked with predicting the pointmap for the second image, explicitly conditioned on the features of the first view. Specifically, for each pixel in the second image, the predicted pointmap value corresponds to the 3D coordinate of its matching pixel in the first image.
The proposed pointmap matching representation facilitates the establishment of explicit 3D point correspondences, which can be directly utilized for motion analysis. Moreover, we further extend our POMATO to 4D video sequence input by introducing a temporal motion module to enhance the learning of temporal motions which helps to guarantee the scale consistency across different frames and improves performance in tasks where both accurate geometry and reliable matching are paramount, most notably 3D point tracking.
The proposed pointmap-based matching representation enables explicit 3D point correspondences, which can be directly utilized for motion analysis. To further extend our approach, we introduce POMATO for 4D video sequence input by incorporating a temporal motion module. This module enhances the learning of temporal motions with scale consistency across frames, and improves performance on tasks where both accurate geometry and robust matching are essential—most notably, 3D point tracking. Compared with the recent temporal 3D reconstruction methods [41, 39] based on an autoregression manner where the previous frames are blocked from the lately added frames, our temporal motion module is based on the self-attention mechanism along the temporal dimension, facilitating a thorough interaction across all frames. Our POMATO is trained in a two-stage manner. In the first stage, we use pairwise input images to learn fundamental geometry and matching capacity. Then we extend the input to sequential video input and insert the temporal motion module, enabling the model to effectively learn motions along the temporal dimension.
Our contributions can be summarized as three folds: First, we propose a novel approach that unifies the fundamental geometry estimation and motion understanding for dynamic 3D reconstruction into a single network by incorporating the representation of pointmap matching. Second, we introduce a temporal motion module to facilitate the interactions of motion features along the temporal dimension which significantly improves the performance in tasks where both accurate geometry and precise matching are required for video sequential input—most notably, 3D point tracking. Third, we demonstrate promising performance on a range of 3D vision tasks, including video depth estimation on dynamic scenes, 3D point tracking, and camera pose estimation.
2 Related Work
Geometry estimation refers to the process of determining the spatial properties and structures from different forms of visual data. Direct recovery of 3D geometry from a single RGB image is by nature an ill-posed problem. Many recent works [21, 49, 3, 14] have tried to leverage strong pre-trained models to learn generalizable depthmaps from large-scale real and synthetic datasets to solve ambiguities. For example, Marigold [21], Geowizard [10], and GenPercept [46] aim at leveraging the generative priors from pre-trained diffusion models by finetuning them on synthetic datasets. Depthanything V2 [49] proposes to estimate scale-and-shift invariant disparity map by finetuning DINOV2 [27] model on synthetic datasets and large-scale pseudo labels. Depth Pro [3] further propose a FOV head to estimate the metric depthmap from a single image without relying on camera intrinsics as input. Due to the scale ambiguity in the monocular depth estimation models, ChronoDepth [34], DepthCrafter [15], and Depth-any-video [48] proposes to learn temporal consistent depthmaps by leveraging the priors from a video generative model, i.e. SVD [2]. In another line of the research, multi-view stereo reconstruction (MVS) methods seek to reconstruct visible surfaces from multiple viewpoints. Traditional MVS [11] and SfM pipelines break the reconstruction pipeline into several sub-problems, e.g., feature extraction [7], image matching [1, 25], triangulation, and bundle adjustment [6]. The chain is complicated and accumulates noise for every single step, thus often resulting in unsatisfactory performance in complex real-world scenes. Recognizing the limitations of previous MVS methods, seminal work DUSt3R [42] proposes 3D pointmaps representation, and trains a network from large-scale data to regress the dense and accurate pointmaps from a pair of images. The camera intrinsics and relative camera poses can be implicitly inferred from the two-view pointmaps. However, it still can not handle reconstruction for dynamic scenes.
Motion representation. Optical flow is a commonly used representation for 2D motion, which is defined as a 2D vector field describing the apparent movements of each pixel between a pair of images. RAFT [36] is a representative work for pairwise optical flow estimation, which employs a 4D cost volume and recurrently estimates the optical flow. Some follow-up methods further extend it to multi-frame (3-5 frames) settings, which is still insufficient for long-range tracking. To resolve the problem, Particle Video [33] represent video motion by using a set of particles. Each particle is an image point sample with a long-duration trajectory and other properties. Particle videos have two key advantages over optical flow: (1) persistence through occlusions, and (2) multi-frame temporal context. Some recent works, PIPs [13], TAPIR [8] and Cotracker [20] have renewed interest in this representation and show promising long-term 2D point tracking results. Recognizing the advantage of point representation, SpatialTracker [45] lifts the 2D points into 3D and performs tracking in the 3D space. Though it can handle occlusions and enhance 3D tracking accuracy, it still relies on a separate monocular depth estimator, which prevents it performing 3D point tracking in an end-to-end fashion.
Multi-view dynamic reconstruction. Our work is closely connected to multi-view dynamic 3D reconstruction techniques. Early works [32, 30] take the straightforward idea that first pre-segment the scene into different regions, each corresponding to a single rigid part of an object, then apply the rigid-SfM technique to each of the regions. However, in a general dynamic setting, the task of densely segmenting rigidly moving objects or parts is not trivial. Some of the recent Neural Radiance Fields (NeRF)  [26] and Gaussian Splatting  [22] based methods have achieved state-of-the-art results. However, most of these methods require simultaneous multi-view video inputs or require predefined templates [16]. Shape of motion [40], proposes a new dynamic scene representation to represent the dynamic scene as a set of persistent 3D Gaussians, and optimize the representation from a monocular video by leveraging monocular depth estimation priors and 2D track estimates across frames. MonST3R directly finetuned the original DUSt3R model upon some synthetic datasets that contain dynamic scenes. Our POMATO also represents the scene as the 3D pointmap, Different from MonST3R [50], we propose a point matching head to unify geometry and matching estimation for dynamic reconstruction.
3 Method
Figure 3: Overview of our training pipeline. (1) Stage I: build upon DUSt3R [42] architecture, we introduce a third regression point-matching head: Head3, which is in parallel to Head2 for explicit pointmap matching in 3D space. For each pixel in the second view, the output pointmap coordinate is the 3D point map of the corresponding pixel in the first view. (2) Stage II: we introduce a temporal fusion module in three heads that enables multi-style sequential input for learning temporal motions.
Figure 4: Inference pipelines for point tracking, video depth, and multi-view reconstruction. $t_{k}$ indicates the keyframe in the sequence.
Figure 5: Architecture of our temporal motion module. We insert a transformer-based motion module into the vanilla DPT [31] head at different feature scales to enhance the consistency along the temporal dimension.
3.1 Preliminary
The overview of our POMATO is demonstrated in Fig.3. We inherit the definition of pointmap ${\bf X}\in{\mathbb{R}}^{H\times W\times 3}$ in DUSt3R [42] as a dense 2D field of 3D points which maps to its corresponding RGB pixels. Given a pair of input images ${\bf I}^{1},{\bf I}^{2}\in{\mathbb{R}}^{H\times W\times 3}$ from two different views, a weight-sharing ViT first extracts the corresponding features ${\bf F}^{1},{\bf F}^{2}$ for each view. Two parallel branches are then introduced to decode the geometric structures and enhance the feature alignment via cross-attention in decoder modules, following a regression head to estimate pointmaps ${\bf X}^{1,1},{\bf X}^{2,1}\in{\mathbb{R}}^{H\times W\times 3}$ along with a confidence map ${\bf C}^{1,1},{\bf C}^{2,1}\in{\mathbb{R}}^{H\times W}$ for each image view. Generally, ${\bf X}^{n,m}$ indicates the pointmap ${\bf X}^{n}$ from camera $n$ expressed in camera $m$’s coordinate frame, which is obtained by a rigid transformation:
${\bf X}^{n,m}={\bf P}_{m}{\bf P}_{n}^{-1}h({\bf X}^{n}),$ (1)
where ${\bf P}_{m},{\bf P}_{n}\in{\mathbb{R}}^{3\times 4}$ are world-to-camera poses for camera $m$ and camera $n$, respectively, and $h({\bf X}^{n})$ is a homogeneous mapping for the 3D coordinate in camera coordinate of camera $n$.
The task for Decoder 1 and its regression head can be briefly summarized as estimating the 3D points for ${\bf I}^{1}$ in its own coordinate system while Decoder 2 and its regression head are responsible for estimating pixel-wise 3D coordinates for ${\bf I}^{2}$ in ${\bf I}^{1}$’s coordinate system after a rigid transformation of global rotation and translation. In the following contents, we will first introduce our POMATO with pairwise input images and then extend it to the video sequence input with our temporal motion module.
3.2 Pointmap Matching with Pairwise Input
As discussed before, the definition of ${\bf X}^{2,1}$ depicts a rigid camera transformation that is ambiguous to reflect explicit matching relationships for dynamic regions. To tackle this, we propose to formulate an explicit pointmap matching ${\bf X}^{2,1}_{m}\in{\mathbb{R}}^{H\times W\times 3}$ that maps dense RGB pixels of ${\bf I}^{2}$ to 3D coordinates of corresponding pixels in ${\bf I}^{1}$ under the first image’s coordinate system. Given a 2D query pixel at $(x_{2},y_{2})$ in ${\bf I}^{2}$ and its corresponding pixel at $(x_{1},y_{1})$ in ${\bf I}^{1}$, the matched pointmap at $(x_{2},y_{2})$ in ${\bf I}^{2}$ is:
${\bf X}^{2,1}_{m}(x_{2},y_{2})={\bf X}^{1,1}(x_{1},y_{1}),$ (2)
where $(x,y)$ indicates the coordinates of 2D grid. For the representative dynamic point (red) in Fig. 2, the pointmap matching result is the 3D coordinate of point A in the coordinate system of the first image. As shown in Fig. 3, ${\bf X}^{2,1}_{m}$ and ${\bf X}^{1,1}$ are supposed to match perfectly in 3D space on the premise of neglecting occluded regions. We argue that the set of decoder tokens from the second branch preserves abundant matching information with iterative cross-attentions, so we introduce a matching head with the same architecture of Head1 and Head2. The supervision for pointmap matching ${\bf X}^{2,1}_{m}$ still follows the 3D regression loss which is defined as the Euclidean distance:
$\mathcal{L}_{\text{m}}=\left\|\frac{1}{z}{{\bf X}^{2,1}_{m}}-\frac{1}{\bar{z}}{\bar{{\bf X}}^{2,1}}_{m}\right\|,\vspace{-1mm}$ (3)
where $\bar{{\bf X}}^{2,1}_{m}$ is the ground truth pointmap matching which can be obtained following Eq. 2 on 2D tracking dataset along with the depth and camera information. $z,\bar{z}$ are the same norm factor defined in DUSt3R. The matching confidence ${\bf C}^{2,1}_{m}$ is also jointly learned with the confidence loss for Head1 and Head2 within valid regions.
$\mathcal{L}_{\text{mconf}}={\bf C}^{2,1}_{m}\mathcal{L}_{\text{m}}-\alpha\text{log}{\bf C}^{2,1}_{m}\vspace{-1mm}$ (4)
The final loss $\mathcal{L}$ of our POMATO for pairwise input is a combination of predefined DUSt3R loss $\mathcal{L}_{\text{DUSt3R}}$, matching loss $\mathcal{L}_{\text{m}}$, and matching confidence loss $\mathcal{L}_{\text{mconf}}$. When training our POMATO for pairwise input images at the first stage, the parameters in the encoder are frozen.
3.3 Dynamic Mask Estimation
Taking advantage of the explicit pointmap matching head, our POMATO can directly perform dynamic mask estimation without introducing a third assistant module like the optical flow model, getting rid of the additional computation cost and the potential domain gap. For an image pair $\{{\bf I}^{i},{\bf I}^{j}\}$ along with the estimation of ${\bf X}^{j,i}$ from Head2 and ${\bf X}^{j,i}_{m}$ from Head3, the dynamic mask ${\bf D}^{j,i}$ can be obtained by comparing the difference between ${\bf X}^{j,i}$ and ${\bf X}^{j,i}_{m}$:
${\bf D}^{j,i}=||{\bf X}^{j,i}_{m}-{\bf X}^{j,i}||>\alpha,$ (5)
where $\alpha$ is a dynamic threshold defined as $3\times\text{median}(||{\bf X}^{j,i}_{m}-{\bf X}^{j,i}||)$. The explicit dynamic mask can be incorporated into the global alignment process to minimize the interference of moving objects for pose estimation and 3D reconstruction. The incorporation of dynamic masks for improving global alignment is detailed in the supplementary materials.
3.4 Motion Module for Video Pointmap Matching
With the fundamental capacity of geometric estimation and pointmap matching for pairwise images, we extend our POMATO to 4D video sequences for temporal-related tasks by inserting a transformer-based motion module into the vanilla DPT [31] head and we refer this enhanced regression head as the ”temporal DPT head”. The architecture of the temporal DPT head is illustrated in Fig. 5. For a set of decoder tokens ${\bf G}\in{\mathbb{R}}^{B,T,N,C}$ where $B,T,N,C$ represent the batch size, window length of a video sequence, token number, and token dimension, respectively, we merge the token number dimension into the batch axis and apply the motion module which consists of two blocks of standard multi-head self-attention modules and feed-forward networks along the temporal dimension $T$. As shown in Fig. 5, the motion module is only applied to decoder tokens of low feature resolutions in terms of minimizing the GPU memory consumption. In this work, the window length $T$ for a video sequence is set to 12 as default. Explorations of different window lengths can be found in Sec. 4. When training with the motion module, we freeze parameters in both the encoder and decoder and finetune the temporal DPT head only. With the introduction of the temporal DPT head, our POMATO can be introduced to applications on the dynamic videos.
3D Point Tracking. Given a video sequence of $T$ frames $\{{\bf I}^{t_{1}},{\bf I}^{t_{2}},...,{\bf I}^{t_{T}}\}$, we can create a set of stereo image pairs: $\{({\bf I}^{t_{k}},{\bf I}^{t_{2}}),({\bf I}^{t_{k}},{\bf I}^{t_{3}}),...,({\bf I}^{t_{k}},{\bf I}^{t_{T}})\}$, where ${\bf I}^{t_{k}}$ is the keyframe. For each pair $({\bf I}^{t_{k}},{\bf I}^{t_{n}})$, where $n\in\{1,2,...T\}$, we compute the dense pointmap matching ${\bf X}^{t_{k},t_{n}}_{m}$ for each reference pixel in the coordinate of the keyframe. As shown in the top part of Fig.4, all the target frames are fed into the Head1 for estimating the pointmap under their own coordinate system while the grouped keyframes are fed to the Head3 for finding the corresponding points in each target frame. The introduced motion module facilitates the interaction along the temporal dimension, enhancing the scale consistency for the global pointmap matching results. Once all pairwise pointmap matching is completed (which can be parallelized using batch operations), we obtain a set of dense 3D tracking results in terms of the keyframe: $\{{\bf X}^{t_{k},t_{2}}_{m},{\bf X}^{t_{k},t_{3}}_{m},...{\bf X}^{t_{k},t_{T}}_{m}\}$. For sparse 3D tracking, the dense tracking results can be sparsified by indexing with the 2D coordinates of query points in ${\bf I}^{t_{k}}$. When tracking across a video sequence longer than $T$ frames, we use a simple sliding-window approach with an overlap of 4 frames for a window length of 12 frames. For the new appearing frames, we directly compute the tracking to the target keyframe. For the overlapped frames, we use a linear weighted method to fuse the prediction from the last sequence to the current. Besides the default loss function for Head1 and Head3, we also introduce a temporal consistency loss by computing the scale factor along the temporal dimension. For pointmap matching loss in Eq.4, the additional temporal tracking loss $\mathcal{L}_{\text{t}}$ is:
$\mathcal{L}_{\text{t}}=\frac{1}{T}\sum^{T}_{i=1}\left\|\frac{1}{z_{T}}{{\bf X}^{t_{k},t_{i}}_{m}}-\frac{1}{\bar{z}_{T}}{\bar{{\bf X}}^{t_{k},t_{i}}}_{m}\right\|,\vspace{-1mm}$ (6)
where the scaling factors $z_{T}=$ norm $({\bf X}^{t_{k},t_{1}}_{m},...,{\bf X}^{t_{k},t_{T}}_{m})$ and $\bar{z}_{T}=$ norm $(\bar{{\bf X}}^{t_{k},t_{1}}_{m},...,\bar{{\bf X}}^{t_{k},t_{T}}_{m})$. The additional temporal loss is similarly applied to the Head1.
Video Depth Estimation. As shown in the middle part of the Fig. 4, the input video sequence is formulated to a set of identical image pairs $\{({\bf I}^{t_{1}},{\bf I}^{t_{1}}),({\bf I}^{t_{2}},{\bf I}^{t_{2}}),...,({\bf I}^{t_{T}},{\bf I}^{t_{T}})\}$. The video depth estimation task involves both Head1 and Head2 where the predictions from each head are identical. We use the output of Head1 as our final depth estimation. Similarly, the temporal consistency loss is also applied to both heads.
3D Reconstruction.Within the window of a video sequence, our POMATO can leverage the advantage of the motion module and perform 3D reconstruction in a feed-forward manner, which skips redundant post-process operations like global alignment. As shown in the bottom part of the Fig. 4, the input video sequence is formulated as $\{({\bf I}^{t_{k}},{\bf I}^{t_{1}}),({\bf I}^{t_{k}},{\bf I}^{t_{2}}),...,({\bf I}^{t_{k}},{\bf I}^{t_{T}})\}$, where ${\bf I}^{t_{k}}$ is the keyframe. The basic idea is to align the pointmap of all reference frames to the coordinate system of the keyframe, and thus all the reference frames are fed to the Head2 while the keyframe is fed to the Head1. With the advantage of the motion module, the pointmap of each reference frame maintains a consistent scale under the same coordinate of the keyframe. Temporal consistency loss, as mentioned in Eq6, is similarly required for both involved heads.
Sintel [4] BONN [28] KITTI [12]
Alignment Method Optim. Onl. Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$ <$1.25\uparrow$
Per-sequence scale DUSt3R-GA [42] 0.656 45.2 0.155 83.3 0.144 81.3
MASt3R-GA [24] 0.641 43.9 0.252 70.1 0.183 74.5
MonST3R-GA [50] 0.378 55.8 0.067 96.3 0.168 74.4
Spann3R [39] 0.622 42.6 0.144 81.3 0.198 73.7
CUT3R [41] 0.421 47.9 0.078 93.7 0.118 88.1
POMATO 0.416 53.6 0.074 96.1 0.085 93.3
Per-sequence scale & shift MonST3R-GA [50] 0.335 58.5 0.063 96.4 0.104 89.5
CUT3R [41] 0.466 56.2 0.111 88.3 0.075 94.3
POMATO 0.345 57.9 0.072 96.5 0.084 93.4
Table 1: Video depth evaluation. We report scale-invariant depth and scale &\&& shift invariant depth accuracy on Sintel [4], Bonn [28], and KITTI [12] datasets. Methods requiring global alignment are marked “GA”, while “Optim.” and “Onl.” indicate optimization-based and online methods, respectively. The best and second best results in each category are bold and underlined, respectively.
4 Experiments
4.1 Experimental Details
Training data. We train our network with a mixture of five datasets: PointOdyssey [52], Tartanair [43], ParallelDomain4D [38], DynamicReplica [19] and Carla (0.9.15) [9]. The specific number and the usage ratio of each dataset can be found in the supplementary materials. All of these datasets are synthetic and come with pixel-accurate ground truth depth, as well as camera intrinsics and extrinsics. They feature diverse dynamic scene types: indoor and outdoor. Among them, PointOdyssey and DynamicReplica have additional 2D trajectory annotations which can be used to construct pointmap matching ground truth following Eq. 2. Thus, all datasets contribute to geometry supervision for training pointmaps in Head1 and Head2 and we use only PointOdyssey, DynamicReplica, and TartanAir datasets to train our proposed pointmap matching head. The training of POMATO follows a two-stage process. In the first stage, we utilize pair-wise data to establish the fundamental capabilities of both geometry and matching. All parameters in the decoders and each DPT head are learnable. In the second stage, we introduce the temporal motion module and freeze parameters in the decoder. The same datasets are used in both stages.
Training and inference details. Our model architecture is based on the publicly available DUSt3R [50] model, utilizing the same backbone consisting of a ViT-Large encoder and a ViT-Base decoder. To maximize the benefits of MonST3R’s geometry estimation ability in dynamic scenes, we initialize the model weights using the publicly available MonST3R checkpoint. For the newly introduced pointmap matching head, we initialize the head weights from the Head2 weight of the MonST3R. We train our network for 10 epochs with a cosine schedule and the initial learning rate is set to 0.0001. The batch size for the first stage of pairwise image training is 16 on 4 A100 (40G) GPUs. When training with the temporal motion module, the batch size is set to 4 with a temporal window length of 12.
4.2 Video Depth Estimation
Following MonST3R [50] and CUT3R [41], we rescale all predictions from the same video to align them together by conducting two forms of alignment: per-sequence scale and shift alignment and per-sequence scale alignment. Thus, we can measure the per-frame depth quality and inter-frame depth consistency. We employ our proposed motion module for video depth estimation as described in Sec.3.4 and compare our method against several variants of DUSt3R, including DUSt3R [42], MAST3R [24], MonST3R [50], Spann3R [39], and CUT3R [41]. Some of these methods [42, 24, 50] rely on global alignment (GA), which assumes a static environment. While GA benefits multi-view consistency in static regions, it can potentially degrade the representation of dynamic elements and is computationally expensive. In contrast, POMATO can perform online video depth inference with the introduced temporal fusion module and achieve much faster inference speed. As shown in Tab. 1, our method demonstrates comparable performance to the global alignment (GA)-based MonST3R [50] on the Sintel [4] and BONN [28] datasets, while surpassing it on KITTI dataset. Besides, we consistently outperform the state-of-the-art online method, CUT3R [41], across various settings. These results underscore the effectiveness of our approach, specifically (1) the joint learning of geometry and pointmap matching, and (2) the temporal motion module.
PointOdyssey [52] ADT [29] PStudio [18] Average
Method T-12 T-24 T-12 T-24 T-12 T-24 T-12 T-24
SpatialTracker [45] 21.24 23.33 22.34 22.87 32.94 32.81 25.51 26.34
DUSt3R [42] 20.27 21.01 30.77 28.64 10.91 6.39 20.65 18.68
MASt3R [24] 17.36 18.29 27.18 25.51 12.83 7.73 19.12 17.17
MonST3R [50] 28.50 29.29 29.82 28.42 18.23 10.81 25.52 22.84
POMATO 34.19 35.75 33.41 30.72 27.05 23.81 31.55 30.09
Table 2: 3D tracking evaluation. We report the APD metric to evaluate 3D point tracking on the PointOdyssey [52], ADT [29], and PStudio [18] datasets. T-12 and T-24 indicate tracking within the temporal length of 12 frames and 24 frames, respectively. POMATO achieves remarkable performance even compared with the specialized model SpatialTracker [45]. * indicates the camera intrinsic is required.
Figure 6: Qualitative comparison of dynamic scenes. Compared to MonST3R, our POMATO can provide more reliable motion masks, 3D point tracking, and reconstruction performance.
4.3 3D Point Tracking
For 3D point tracking task, we use the Aria Digital Twin (ADT) [29], and Panoptic Studio (PStudio) [18] benchmarks from the TAPVid-3D [23] dataset along with the validation set on the PointOdyssey [52] dataset. We report the Average Percent Deviation (APD) metric, which quantifies the average percentage of points within a threshold relative to the ground truth depth. The APD metric serves as a direct measure of the accuracy of the predicted tracking. We reformulate the datasets and project all the query points within a testing sequence to the first frame. We report tracking results on a length of both 12 and 24 frames. As shown in Tab.2, our POMATO achieves the best performance on both PointOdyssey and ADT datasets. Additionally, our method demonstrates superior generalization, as reflected by the highest average APD metric. It’s worth mentioning that SpatialTracker [45] is a state-of-the-art network tailored for 3D point tracking with ground truth camera intrinsic as additional input data. POMATO surpasses it on two datasets and improves the average APD metric by 23.7% and 14.2% for 12 frames and 24 frames, respectively. In contrast, DuST3R-based methods struggle with ambiguous matching representations, resulting in imprecise tracking performance in dynamic scenarios.
Video Depth Tracking
Temporal Length Sintel [4] Bonn [28] KITTI [12] ADT [29] PStudio [18] PointOdyssey [52]
[…]
Pair-wise 0.548 46.2 0.087 94.0 0.113 89.5 31.35 25.20 33.21
6 frames 0.436 51.3 0.076 95.9 0.085 93.5 32.63 26.93 33.88
12 frames 0.416 53.6 0.075 96.1 0.086 93.3 33.41 27.05 34.19
Table 3: Ablation study on the temporal motion module. The introduction of the temporal motion module brings a significant improvement compared with the fundamental model trained on pairwise images. As the temporal window length enlarges from 6 frames to 12 frames, we obtain an overall consistent improvement for both video depth estimation and 3D point tracking tasks.
TUM [35] Bonn [28]
Method ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$ ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$
DUSt3R [42] 0.025 0.013 2.361 0.030 0.025 2.522
MASt3R [24] 0.027 0.015 1.910 0.031 0.025 2.478
MonST3R [50] 0.021 0.006 1.142 0.025 0.021 2.120
CUT3R [41] 0.023 0.016 0.510 0.028 0.033 2.569
POMATO 0.020 0.010 0.509 0.037 0.016 1.837
Table 4: Pose estimation. * indicates using an off-the-shelf optical flow model to get the motion mask and the pseudo 2D optical flow ground truth. Our method achieves an overall best performance and improves the RPE rot metric significantly.
Figure 7: Effectiveness of our motion mask estimation in 3D reconstruction. Without explicitly filtering out the motion area, both pose and geometry estimation will be degenerated.
4.4 Camera Pose Estimation
We conduct pose estimation experiments on Bonn  [28] and TUM [35] datasets which include dynamic moving objects, and compare our method with the DUSt3R-related works along with the recently proposed CUT3R[41]. Global alignment is utilized for the evaluation. The sampling stride is set to 5 on the Bonn dataset and 3 on the TUM dataset. We evaluate over 40 frames and report the results in Tab. 4. Three metrics are reported: Absolute Translation Error (ATE), Relative Translation Error (RPE trans), and Relative Rotation Error (RPE rot). Our method achieves the best overall performance. Notably, POMATO operates without relying on any auxiliary modules for dynamic object estimation, thereby avoiding potential domain gaps across modules and preventing error accumulation. In particular, POMATO significantly improves the RPE-rot metric, surpassing MonST3R by 55.4% and 13.3% on the TUM and Bonn datasets, respectively. We present a visualization to demonstrate the effectiveness of our dynamic mask estimation in 3D reconstruction, as mentioned in Sec.3.3. Without explicitly identifying dynamic regions, both camera pose and geometry estimation suffer from significant degeneration.
4.5 Ablation Study
Extensive ablation experiments are conducted on video depth estimation and 3D point tracking tasks to validate the effectiveness of learning temporal motions on video sequence input. We report three models trained with pairwise images, a shorter temporal window length of 6 frames, and the default temporal window length of 12 frames. As shown in Tab. 3, the introduction of the temporal motion module introduces significant improvements across all datasets, emphasizing the importance of consistency among different frames. When enlarging the temporal window length from 6 frames to 12 frames, the video depth estimation obtains a further improvement on both Sintel and Bonn datasets. For 3D point tracking, the enlarged temporal window length brings a consistent enhancement across all three testing datasets.
5 Discussion and Conclusion
We introduce POMATO, a unified framework designed for geometry estimation and motion understanding in dynamic scenes. Leveraging our proposed pointmap matching head, our method can effectively distinguish moving regions, thereby mitigating the interference caused by dynamic objects. The introduced temporal motion module enhances the learning of temporal motions across different frames, improving the scale consistency and boosting the performance in tasks where geometry and matching are critical, most notably, 3D point tracking. Moving forward, we aim to explore methods to scale up our training with more matching data and further improve the 3D reconstruction performance.
Supplementary Material
Dataset Domain Scene Type # of Frames # of Scenes Dynamics Ratio
PointOdyssey [52] Synthetic Indoors & Outdoors 200k 131 Realistic 57.1%
TartanAir [43] Synthetic Indoors & Outdoors 100k 163 None 14.3%
DynamicReplica [19] Synthetic Indoors 145k 524 Realistic 14.3%
ParallelDomain4D [38] Synthetic Outdoors 750k 15015 Driving 8.6%
Carla [9] Synthetic Outdoors 7k 5 Driving 5.7%
Table 5: An overview of all training datasets and sample ratio. All datasets provide both camera pose, depth, and most of them include dynamic objects.
Figure 8: Fast 3D reconstruction with our temporal motion module. Given a sequence of images less than temporal window length, our POMATO can directly obtain a global pointmap under the key frame coordinate.
Appendix A Pointmap Matching for Global Alignment.
Given a sequence of video frames, the target of global alignment is to project all pairwise estimated pointmaps to the same global world coordinates. DUSt3R constructs a connectivity pairwise graph and aims to minimize the re-projection error for each image pair globally where the dynamic regions are supposed to be separated from the static regions. To this end, MonST3R [50] further introduces an assistant optical flow network [44] to help mask the dynamic regions and provide a pseudo label of 2D matching for minimizing the re-projection error in static regions. However, the introduced assistant model will introduce inevitable domain gaps and additional computation costs. Besides, the optical flow model is tailored for matching within two adjacent frames, suffering an obvious degeneration with the large view displacement. In POMATO, for an image pair $\{{\bf I}^{i},{\bf I}^{j}\}$, the dynamic mask ${\bf D}^{j,i}$ is calculated by comparing the difference between ${\bf X}^{j,i}$ and ${\bf X}^{j,i}_{m}$:
${\bf D}^{j,i}=||{\bf X}^{j,i}_{m}-{\bf X}^{j,i}||>\alpha,$ (7)
where $\alpha$ is a dynamic threshold defined as $3\times\text{median}(||{\bf X}^{j,i}_{m}-{\bf X}^{j,i}||)$.
Given the updated camera intrinsic $\tilde{K}$ after an iteration of optimization, the target matching 2D coordinates ${\bf F}^{j,i}_{m}\in{\mathbb{R}}^{H\times W\times 2}$ can be calculated as ${\bf F}^{j,i}_{m}=p(\tilde{\bf K}{\bf X}^{j,i}_{m})$ where $p$ is a mapping from 3D camera coordinates to 2D pixel coordinates. The optical flow loss proposed in MonST3R can thus be modified with our dynamic mask and 2D matching coordinates. Details about the optical flow loss are referred to MonST3R [50].
Appendix B Fast 3D Reconstruction with video POMATO
Given a sequence of images less than the temporal window length of 12 frames, dynamic 3D reconstruction can be obtained by directly estimating the pointmaps of all reference images to the coordinate of the key frame as discussed in the Sec.3.4. Here, we provide more visualization results of this feed-forward manner and demonstrate the effectiveness of introducing the temporal motion module. As shown in Fig.8, directly applying the pairwise reconstruction will suffer from an obvious scale shift among different frames. After the temporal motion module, the consistency within the video sequence obtains an obvious enhancement.
Appendix C Training Data Details
The details about the training datasets can be found in Tab.5. The finetuning procedure of POMATO was conducted exclusively using synthetic training datasets.
Appendix D More Visualizations on Dynamic Scenes
We provide more visualizations in Fig. 9 and Fig. 10. MonST3R suffers obvious degeneration when the view displacement is large as reflected by the erroneous pose estimation while POMATO can still provide a consistent camera trajectory.
Figure 9: Compared with MonST3R, our POMATO can provide more complete dynamic masks and consistent geometry.
Figure 10: MonST3R suffers obvious degeneration when the view displacement is large as reflected by the erroneous pose estimation while POMATO can still provide a consistent camera trajectory.

### v2

@@
 Abstract
-3D reconstruction in dynamic scenes primarily relies on the combination of geometry estimation and matching modules where the latter task is pivotal for distinguishing dynamic regions which can help to mitigate the interference introduced by camera and object motion.
+Recent approaches to 3D reconstruction in dynamic scenes primarily rely on the integration of separate geometry estimation and matching modules, where the latter plays a critical role in distinguishing dynamic regions and mitigating the interference caused by moving objects.
 Furthermore, the matching module explicitly models object motion, enabling the tracking of specific targets and advancing motion understanding in complex scenarios.
-Recently, the proposed representation of pointmap in DUSt3R suggests a potential solution to unify both geometry estimation and matching in 3D space, but it still struggles with ambiguous matching in dynamic regions, which may hamper further improvement.
+Recently, the proposed representation of pointmap in DUSt3R suggests a potential solution to unify both geometry estimation and matching in 3D space, effectively reducing computational overhead by eliminating the need for redundant auxiliary modules.
+However, it still struggles with ambiguous correspondences in dynamic regions, which limits reconstruction performance in such scenarios.
 In this work, we present POMATO, a unified framework for dynamic 3D reconstruction by marrying POintmap MAtching with Temporal mOtion.
-Specifically, our method first learns an explicit matching relationship by mapping RGB pixels from both dynamic and static regions across different views to 3D pointmaps within a unified coordinate system.
-Furthermore, we introduce a temporal motion module for dynamic motions that ensures scale consistency across different frames and enhances performance in tasks requiring both precise geometry and reliable matching, most notably 3D point tracking.
-We show the effectiveness of the proposed pointmap matching and temporal fusion paradigm by demonstrating the remarkable performance across multiple downstream tasks, including video depth estimation, 3D point tracking, and pose estimation.
+Specifically, our method first learns an explicit matching relationship by mapping RGB pixels across different views to 3D pointmaps within a unified coordinate system.
+Furthermore, we introduce a temporal motion module for dynamic motions that ensures scale consistency across different frames and enhances performance in 3D reconstruction tasks requiring both precise geometry and reliable matching, most notably 3D point tracking.
+We show the effectiveness of our proposed POMATO by demonstrating the remarkable performance across multiple downstream tasks, including video depth estimation, 3D point tracking, and pose estimation.
 Code and models are publicly available at https://github.com/wyddmw/POMATO.
@@
 Image-based 3D reconstruction is a fundamental task in computer vision with a wide range of applications including SLAM [39], robotics [19, 49], autonomous driving [53], and novel view synthesis [5].
-While substantial progress has been achieved in static 3D reconstruction [14, 49, 21, 42, 24], dynamic scenes remain significantly more challenging.
-The presence of moving objects introduces random motion and deformation, which can interfere with the learning of local structure and camera motion, thus complicating accurate geometry estimation.
-These scenarios require joint modeling of both scene geometry and object motion.
-Moreover, downstream tasks such as 3D point tracking demand precise geometry estimation and robust matching across views.
-To effectively distinguish dynamic regions, it is essential to establish reliable correspondences between frames.
-Some pioneering works have attempted to address dynamic motion by incorporating auxiliary matching modules, such as optical flow [40, 50] or 2D tracking [45].
+While substantial progress has been achieved in static 3D reconstruction [16, 51, 23, 44, 26], dynamic scenes remain a major hurdle due to complexities like non-rigid motion and deformation, which may hamper the learning of local structure and camera motion, thereby complicating accurate 3D reconstruction for dynamic scenes.
+These scenarios require explicit modeling of both scene geometry and object motion.
+Moreover, downstream reconstruction tasks, such as 3D point tracking, demand precise geometry estimation and robust matching across views.
+To effectively distinguish dynamic regions, it is essential to establish reliable correspondences between different frames.
+Some pioneering works have attempted to address dynamic motion by incorporating additional auxiliary matching modules, such as optical flow [42, 52] or 2D tracking [47].
 However, these approaches may suffer from domain gaps and accumulated errors between modules, limiting their effectiveness.
@@
 The network utilizes a standard transformer-based encoder-decoder architecture and receives a pair of images as input.
-The system incorporates two parallel decoders with a regression head to predict pointmaps for each view within the same coordinate system as the initial view.
-However, the representation of the pointmap still faces limitations when handling ambiguous correspondence in dynamic scenes, as illustrated in Fig.
+The system incorporates two parallel decoders to predict pointmaps for each view within the same coordinate system.
+However, this representation is limited to static matching and struggles in dynamic scenes, as illustrated in Fig.
 2.
-In static regions, the pointmap of the second image accurately reflects the rigid camera transformation from the first view, where identical RGB pixels correspond to the same 3D coordinates (orange point).
-Conversely, in dynamic regions, pixel displacements or deformations due to object motion introduce inconsistencies in the predicted pointmaps across views (red point), resulting in inaccurate 3D correspondences.
-This inconsistency fails to depict the accurate matching relationships in 3D space.
-To tackle this problem, we present POMATO, a unified network for dynamic 3D reconstruction by marrying POintmap MAtching with Temporal mOtion.
-We introduce an auxiliary matching head that leverages decoder tokens from the second image while preserving matching features through iterative cross-attention across views.
-This module is tasked with predicting the pointmap for the second image, explicitly conditioned on the features of the first view.
-Specifically, for each pixel in the second image, the predicted pointmap value corresponds to the 3D coordinate of its matching pixel in the first image.
-The proposed pointmap matching representation facilitates the establishment of explicit 3D point correspondences, which can be directly utilized for motion analysis.
-Moreover, we further extend our POMATO to 4D video sequence input by introducing a temporal motion module to enhance the learning of temporal motions which helps to guarantee the scale consistency across different frames and improves performance in tasks where both accurate geometry and reliable matching are paramount, most notably 3D point tracking.
-The proposed pointmap-based matching representation enables explicit 3D point correspondences, which can be directly utilized for motion analysis.
-To further extend our approach, we introduce POMATO for 4D video sequence input by incorporating a temporal motion module.
-This module enhances the learning of temporal motions with scale consistency across frames, and improves performance on tasks where both accurate geometry and robust matching are essential—most notably, 3D point tracking.
-Compared with the recent temporal 3D reconstruction methods [41, 39] based on an autoregression manner where the previous frames are blocked from the lately added frames, our temporal motion module is based on the self-attention mechanism along the temporal dimension, facilitating a thorough interaction across all frames.
+To address this problem, we present POMATO, a unified network for dynamic 3D reconstruction by marrying POintmap MAtching with Temporal mOtion.
+We argue that with iterative cross-attention modules across different views, matching features are well preserved in the decoder tokens.
+We thus introduce an auxiliary pointmap matching head to learn explicit correspondences.
+Specifically, for each pixel in the second view, the pointmap matching head predicts the corresponding 3D coordinates of its counterpart in the first view, under the shared coordinate system.
+Our proposed pointmap-based matching representation enables the establishment of explicit correspondences in 3D space, which can be directly leveraged for motion analysis, especially the estimation of dynamic regions.
+Moreover, we further extend our POMATO to handle 4D video sequences by introducing a temporal motion module that enhances the learning of temporal motions.
+This motion module promotes scale consistency across different frames and improves performance in tasks where both accurate geometry and reliable matching are paramount, most notably 3D point tracking.
+Compared with recent temporal 3D reconstruction methods [43, 41] based on an autoregression manner where the previous frames are blocked from the recently added frames, our temporal motion module is based on the self-attention mechanism along the temporal dimension, facilitating a comprehensive interaction across all frames.
 Our POMATO is trained in a two-stage manner.
-In the first stage, we use pairwise input images to learn fundamental geometry and matching capacity.
-Then we extend the input to sequential video input and insert the temporal motion module, enabling the model to effectively learn motions along the temporal dimension.
-Our contributions can be summarized as three folds: First, we propose a novel approach that unifies the fundamental geometry estimation and motion understanding for dynamic 3D reconstruction into a single network by incorporating the representation of pointmap matching.
-Second, we introduce a temporal motion module to facilitate the interactions of motion features along the temporal dimension which significantly improves the performance in tasks where both accurate geometry and precise matching are required for video sequential input—most notably, 3D point tracking.
-Third, we demonstrate promising performance on a range of 3D vision tasks, including video depth estimation on dynamic scenes, 3D point tracking, and camera pose estimation.
+In the first stage, we used pairwise input images to learn fundamental geometry and matching capabilities.
+In the second stage, we extend the input to sequential video frames and incorporate the temporal motion module, enabling the model to effectively capture motions over time.
+Our contributions can be summarized in threefold: First, we propose a novel approach that unifies the fundamental geometry estimation and motion understanding for dynamic 3D reconstruction into a single network by incorporating the representation of pointmap matching.
+Second, we introduce a temporal motion module to facilitate the interactions of motion features along the temporal dimension, which significantly improves the performance in tasks where both accurate geometry and precise matching are required for video sequential input, most notably 3D point tracking.
+Third, we demonstrate promising performance on 3D vision tasks, including video depth estimation, 3D point tracking, and camera pose estimation.
 2 Related Work
@@
 However, it still can not handle reconstruction for dynamic scenes.
+MonST3R [52] directly finetuned the original DUSt3R model upon synthetic datasets that contain dynamic scenes.
 Motion representation.
-Optical flow is a commonly used representation for 2D motion, which is defined as a 2D vector field describing the apparent movements of each pixel between a pair of images.
+Optical flow is a commonly used representation for 2D motion.
 RAFT [38] is a representative work for pairwise optical flow estimation, which employs a 4D cost volume and recurrently estimates the optical flow.
@@
 Early works [34, 32] take the straightforward idea that first pre-segment the scene into different regions, each corresponding to a single rigid part of an object, then apply the rigid-SfM technique to each of the regions.
-However, in a general dynamic setting, the task of densely segmenting rigidly moving objects or parts is not trivial.
 Some of the recent Neural Radiance Fields (NeRF)  [28] and Gaussian Splatting  [24] based methods have achieved state-of-the-art results.
@@
 Shape of motion [42], proposes a new dynamic scene representation to represent the dynamic scene as a set of persistent 3D Gaussians, and optimize the representation from a monocular video by leveraging monocular depth estimation priors and 2D track estimates across frames.
-MonST3R directly finetuned the original DUSt3R model upon some synthetic datasets that contain dynamic scenes.
-Our POMATO also represents the scene as the 3D pointmap, Different from MonST3R [50], we propose a point matching head to unify geometry and matching estimation for dynamic reconstruction.
 3 Method
@@
 (2) Stage II: we introduce a temporal fusion module in three heads that enables multi-style sequential input for learning temporal motions.
-Figure 4: Inference pipelines for point tracking, video depth, and multi-view reconstruction. $t_{k}$ indicates the keyframe in the sequence.
-Figure 5: Architecture of our temporal motion module.
-We insert a transformer-based motion module into the vanilla DPT [31] head at different feature scales to enhance the consistency along the temporal dimension.
 3.1 Preliminary
 The overview of our POMATO is demonstrated in Fig.3.
-We inherit the definition of pointmap ${\bf X}\in{\mathbb{R}}^{H\times W\times 3}$ in DUSt3R [42] as a dense 2D field of 3D points which maps to its corresponding RGB pixels.
+We adopt the definition of pointmap ${\bf X}\in{\mathbb{R}}^{H\times W\times 3}$ in DUSt3R [44] as a dense 2D field of 3D points where each point corresponds to its respective RGB pixel.
 Given a pair of input images ${\bf I}^{1},{\bf I}^{2}\in{\mathbb{R}}^{H\times W\times 3}$ from two different views, a weight-sharing ViT first extracts the corresponding features ${\bf F}^{1},{\bf F}^{2}$ for each view.
-Two parallel branches are then introduced to decode the geometric structures and enhance the feature alignment via cross-attention in decoder modules, following a regression head to estimate pointmaps ${\bf X}^{1,1},{\bf X}^{2,1}\in{\mathbb{R}}^{H\times W\times 3}$ along with a confidence map ${\bf C}^{1,1},{\bf C}^{2,1}\in{\mathbb{R}}^{H\times W}$ for each image view.
+Two parallel branches are employed to decode the geometric structures and enhance the feature alignment via cross-attention in decoder modules, following a regression head to estimate pointmaps ${\bf X}^{1,1},{\bf X}^{2,1}\in{\mathbb{R}}^{H\times W\times 3}$ along with a confidence map ${\bf C}^{1,1},{\bf C}^{2,1}\in{\mathbb{R}}^{H\times W}$ for each image view.
 Generally, ${\bf X}^{n,m}$ indicates the pointmap ${\bf X}^{n}$ from camera $n$ expressed in camera $m$’s coordinate frame, which is obtained by a rigid transformation:
@@
 where ${\bf P}_{m},{\bf P}_{n}\in{\mathbb{R}}^{3\times 4}$ are world-to-camera poses for camera $m$ and camera $n$, respectively, and $h({\bf X}^{n})$ is a homogeneous mapping for the 3D coordinate in camera coordinate of camera $n$.
-The task for Decoder 1 and its regression head can be briefly summarized as estimating the 3D points for ${\bf I}^{1}$ in its own coordinate system while Decoder 2 and its regression head are responsible for estimating pixel-wise 3D coordinates for ${\bf I}^{2}$ in ${\bf I}^{1}$’s coordinate system after a rigid transformation of global rotation and translation.
+The task for Decoder 1 and its regression head estimate the 3D points for ${\bf I}^{1}$ in its own coordinate system while Decoder 2 and its regression head are responsible for estimating pixel-wise 3D coordinates for ${\bf I}^{2}$ in ${\bf I}^{1}$’s coordinate system after a rigid transformation of global rotation and translation.
 In the following contents, we will first introduce our POMATO with pairwise input images and then extend it to the video sequence input with our temporal motion module.
@@
 The supervision for pointmap matching ${\bf X}^{2,1}_{m}$ still follows the 3D regression loss which is defined as the Euclidean distance:
-$\mathcal{L}_{\text{m}}=\left\|\frac{1}{z}{{\bf X}^{2,1}_{m}}-\frac{1}{\bar{z}}{\bar{{\bf X}}^{2,1}}_{m}\right\|,\vspace{-1mm}$ (3)
-where $\bar{{\bf X}}^{2,1}_{m}$ is the ground truth pointmap matching which can be obtained following Eq.
-2 on 2D tracking dataset along with the depth and camera information.
-$z,\bar{z}$ are the same norm factor defined in DUSt3R.
-The matching confidence ${\bf C}^{2,1}_{m}$ is also jointly learned with the confidence loss for Head1 and Head2 within valid regions.
-$\mathcal{L}_{\text{mconf}}={\bf C}^{2,1}_{m}\mathcal{L}_{\text{m}}-\alpha\text{log}{\bf C}^{2,1}_{m}\vspace{-1mm}$ (4)
+$\mathcal{L}_{\text{m}}=\left\|\frac{1}{z_{m}}{{\bf X}^{2,1}_{m}}-\frac{1}{\bar{z}_{m}}{\bar{{\bf X}}^{2,1}}_{m}\right\|,\vskip-2.84526pt$ (3)
+where $\bar{{\bf X}}^{2,1}_{m}$ is the ground truth pointmap matching, which can be obtained following Eq.
+2 on the 2D tracking dataset with the depth and camera information.
+$z_{m},\bar{z}_{m}$ are the same norm factor defined in DUSt3R.
+The matching confidence ${\bf C}^{2,1}_{m}$ is also learned following the confidence loss for Head1 and Head2 within valid regions:
+$\mathcal{L}_{\text{mconf}}={\bf C}^{2,1}_{m}\mathcal{L}_{\text{m}}-\alpha\text{log}{\bf C}^{2,1}_{m}\vskip-2.84526pt$ (4)
 The final loss $\mathcal{L}$ of our POMATO for pairwise input is a combination of predefined DUSt3R loss $\mathcal{L}_{\text{DUSt3R}}$, matching loss $\mathcal{L}_{\text{m}}$, and matching confidence loss $\mathcal{L}_{\text{mconf}}$.
@@
 3.3 Dynamic Mask Estimation
-Taking advantage of the explicit pointmap matching head, our POMATO can directly perform dynamic mask estimation without introducing a third assistant module like the optical flow model, getting rid of the additional computation cost and the potential domain gap.
+Taking advantage of the explicit pointmap matching head, our POMATO can directly perform dynamic mask estimation without introducing an assistant module such as the optical flow model, getting rid of the additional computation cost and the potential domain gap.
 For an image pair $\{{\bf I}^{i},{\bf I}^{j}\}$ along with the estimation of ${\bf X}^{j,i}$ from Head2 and ${\bf X}^{j,i}_{m}$ from Head3, the dynamic mask ${\bf D}^{j,i}$ can be obtained by comparing the difference between ${\bf X}^{j,i}$ and ${\bf X}^{j,i}_{m}$:
-${\bf D}^{j,i}=||{\bf X}^{j,i}_{m}-{\bf X}^{j,i}||>\alpha,$ (5)
-where $\alpha$ is a dynamic threshold defined as $3\times\text{median}(||{\bf X}^{j,i}_{m}-{\bf X}^{j,i}||)$.
-The explicit dynamic mask can be incorporated into the global alignment process to minimize the interference of moving objects for pose estimation and 3D reconstruction.
-The incorporation of dynamic masks for improving global alignment is detailed in the supplementary materials.
-3.4 Motion Module for Video Pointmap Matching
-With the fundamental capacity of geometric estimation and pointmap matching for pairwise images, we extend our POMATO to 4D video sequences for temporal-related tasks by inserting a transformer-based motion module into the vanilla DPT [31] head and we refer this enhanced regression head as the ”temporal DPT head”.
-The architecture of the temporal DPT head is illustrated in Fig.
-5.
+[…]
+[…]
+Figure 4: Architecture of our temporal motion module. We insert a transformer-based motion module (in shallow yellow) into the vanilla DPT [33] head to enhance the temporal consistency.
+3.4 Temporal Motion Module
+With the fundamental capability of geometric estimation and pointmap matching for pairwise images, we follow [6] and extend our POMATO to 4D video sequences by inserting a transformer-based motion module into the vanilla DPT head to construct the ”temporal DPT head”, which is illustrated in Fig.4.
 For a set of decoder tokens ${\bf G}\in{\mathbb{R}}^{B,T,N,C}$ where $B,T,N,C$ represent the batch size, window length of a video sequence, token number, and token dimension, respectively, we merge the token number dimension into the batch axis and apply the motion module which consists of two blocks of standard multi-head self-attention modules and feed-forward networks along the temporal dimension $T$.
-As shown in Fig.
-5, the motion module is only applied to decoder tokens of low feature resolutions in terms of minimizing the GPU memory consumption.
-In this work, the window length $T$ for a video sequence is set to 12 as default.
-Explorations of different window lengths can be found in Sec.
-4.
-When training with the motion module, we freeze parameters in both the encoder and decoder and finetune the temporal DPT head only.
-With the introduction of the temporal DPT head, our POMATO can be introduced to applications on the dynamic videos.
+To reduce the computation cost, the temporal motion modules are applied to features of low resolution.
+3.5 Downstream Temporal Tasks
+Given a video sequence of $T$ frames ${\bf I}^{t_{1}},{\bf I}^{t_{2}},\ldots,{\bf I}^{t_{T}}$}, we construct a unique set of stereo image pairs for each task.
+As illustrated in Fig.
+5, the flexible construction of input pairs—combined with the proposed temporal motion module and pointmap matching head—enables POMATO to seamlessly address downstream temporal tasks, including 3D point tracking, video depth estimation, and 3D reconstruction.
+The keyframe selection strategy and input formulation for each task are detailed in the following section.
+Besides the default regression losses for Head1 and Head2, and predefined losses Eq.
+3 and Eq.
+4 for Head3, we further employ a temporal consistency loss, $\mathcal{L}_{\text{t}}$, which will be described in detail below.
+In addition to the default regression losses for Head1 and Head2, and the predefined losses in Eq.
+3 and Eq.
+4 for Head3, we further introduce a temporal consistency loss, $\mathcal{L}_{\text{t}}$, which will also be described in detail below.
+Figure 5: Inference pipelines for point tracking, video depth, and multi-view reconstruction.
+$t_{k}$ indicates the keyframe.
+With the help of the motion module and flexible input construction, POMATO can be easily applied to downstream temporal tasks.
 3D Point Tracking.
-Given a video sequence of $T$ frames $\{{\bf I}^{t_{1}},{\bf I}^{t_{2}},...,{\bf I}^{t_{T}}\}$, we can create a set of stereo image pairs: $\{({\bf I}^{t_{k}},{\bf I}^{t_{2}}),({\bf I}^{t_{k}},{\bf I}^{t_{3}}),...,({\bf I}^{t_{k}},{\bf I}^{t_{T}})\}$, where ${\bf I}^{t_{k}}$ is the keyframe.
-For each pair $({\bf I}^{t_{k}},{\bf I}^{t_{n}})$, where $n\in\{1,2,...T\}$, we compute the dense pointmap matching ${\bf X}^{t_{k},t_{n}}_{m}$ for each reference pixel in the coordinate of the keyframe.
-As shown in the top part of Fig.4, all the target frames are fed into the Head1 for estimating the pointmap under their own coordinate system while the grouped keyframes are fed to the Head3 for finding the corresponding points in each target frame.
-The introduced motion module facilitates the interaction along the temporal dimension, enhancing the scale consistency for the global pointmap matching results.
-Once all pairwise pointmap matching is completed (which can be parallelized using batch operations), we obtain a set of dense 3D tracking results in terms of the keyframe: $\{{\bf X}^{t_{k},t_{2}}_{m},{\bf X}^{t_{k},t_{3}}_{m},...{\bf X}^{t_{k},t_{T}}_{m}\}$.
-For sparse 3D tracking, the dense tracking results can be sparsified by indexing with the 2D coordinates of query points in ${\bf I}^{t_{k}}$.
-When tracking across a video sequence longer than $T$ frames, we use a simple sliding-window approach with an overlap of 4 frames for a window length of 12 frames.
-For the new appearing frames, we directly compute the tracking to the target keyframe.
-For the overlapped frames, we use a linear weighted method to fuse the prediction from the last sequence to the current.
-Besides the default loss function for Head1 and Head3, we also introduce a temporal consistency loss by computing the scale factor along the temporal dimension.
-For pointmap matching loss in Eq.4, the additional temporal tracking loss $\mathcal{L}_{\text{t}}$ is:
-$\mathcal{L}_{\text{t}}=\frac{1}{T}\sum^{T}_{i=1}\left\|\frac{1}{z_{T}}{{\bf X}^{t_{k},t_{i}}_{m}}-\frac{1}{\bar{z}_{T}}{\bar{{\bf X}}^{t_{k},t_{i}}}_{m}\right\|,\vspace{-1mm}$ (6)
-where the scaling factors $z_{T}=$ norm $({\bf X}^{t_{k},t_{1}}_{m},...,{\bf X}^{t_{k},t_{T}}_{m})$ and $\bar{z}_{T}=$ norm $(\bar{{\bf X}}^{t_{k},t_{1}}_{m},...,\bar{{\bf X}}^{t_{k},t_{T}}_{m})$.
-The additional temporal loss is similarly applied to the Head1.
-Video Depth Estimation.
-As shown in the middle part of the Fig.
-4, the input video sequence is formulated to a set of identical image pairs $\{({\bf I}^{t_{1}},{\bf I}^{t_{1}}),({\bf I}^{t_{2}},{\bf I}^{t_{2}}),...,({\bf I}^{t_{T}},{\bf I}^{t_{T}})\}$.
-The video depth estimation task involves both Head1 and Head2 where the predictions from each head are identical.
-We use the output of Head1 as our final depth estimation.
-Similarly, the temporal consistency loss is also applied to both heads.
-3D Reconstruction.Within the window of a video sequence, our POMATO can leverage the advantage of the motion module and perform 3D reconstruction in a feed-forward manner, which skips redundant post-process operations like global alignment.
-As shown in the bottom part of the Fig.
-4, the input video sequence is formulated as $\{({\bf I}^{t_{k}},{\bf I}^{t_{1}}),({\bf I}^{t_{k}},{\bf I}^{t_{2}}),...,({\bf I}^{t_{k}},{\bf I}^{t_{T}})\}$, where ${\bf I}^{t_{k}}$ is the keyframe.
-The basic idea is to align the pointmap of all reference frames to the coordinate system of the keyframe, and thus all the reference frames are fed to the Head2 while the keyframe is fed to the Head1.
-With the advantage of the motion module, the pointmap of each reference frame maintains a consistent scale under the same coordinate of the keyframe.
-Temporal consistency loss, as mentioned in Eq6, is similarly required for both involved heads.
+As illustrated at the top of Fig.5, the keyframe is set to the first image of the global video sequence and fed to the proposed Head3 to obtain the pointmap matching result of each query point (initialized at the first image) under the coordinate system of each reference frame $\{{\bf X}^{t_{1},t_{1}}_{m},{\bf X}^{t_{1},t_{2}}_{m},{\bf X}^{t_{1},t_{3}}_{m},...{\bf X}^{t_{1},t_{T}}_{m}\}$, while the set of reference frames $\{{\bf I}^{t_{1}},{\bf I}^{t_{2}},{\bf I}^{t_{3}},...{\bf I}^{t_{T}}\}$ is fed to the Head1 to obtain the pointmap under each ego coordinate system.
+The dense tracking results can be further sparsified by indexing the 2D coordinates.
+When inference on a video longer than $T$ frames, a simple sliding-window approach with an overlap of four frames is adopted to enhance the consistency between adjacent video windows.
+The temporal consistency loss $\mathcal{L}_{\text{t}}$ for tracking is:
+$\mathcal{L}_{\text{t}}=\frac{1}{T}\sum_{i=1}^{T}\left\|\frac{{{\bf X}^{t_{1},t_{i}}_{m}}}{z_{m}^{T}}-\frac{{\bar{{\bf X}}^{t_{1},t_{i}}}_{m}}{\bar{z}_{m}^{T}}\right\|+\left\|\frac{{{\bf X}^{t_{i},t_{i}}}}{z^{T}}-\frac{{\bar{{\bf X}}^{t_{i},t_{i}}}}{\bar{z}^{T}}\right\|,$ (6)
+where $z_{m}^{T}=$ norm $({\bf X}^{t_{1},t_{1}}_{m},{\bf X}^{t_{1},t_{2}}_{m},...,{\bf X}^{t_{1},t_{T}}_{m})$ and $\bar{z}_{T}=$ norm $(\bar{{\bf X}}^{t_{1},t_{1}}_{m},\bar{{\bf X}}^{t_{1},t_{2}}_{m},...,\bar{{\bf X}}^{t_{1},t_{T}}_{m})$.
+$z_{m}^{T}$ and $\bar{z}_{T}$ are defined analogously.
+[…]
+$\mathcal{L}_{\text{t}}=\frac{1}{T}\sum_{i=1}^{T}\left\|\frac{{{\bf X}^{t_{i},t_{i}}_{1}}}{z_{1}^{T}}-\frac{{\bar{{\bf X}}^{t_{i},t_{i}}}}{\bar{z}^{T}}\right\|+\left\|\frac{{{\bf X}_{2}^{t_{i},t_{i}}}}{z_{2}^{T}}-\frac{{\bar{{\bf X}}^{t_{i},t_{i}}}}{\bar{z}^{T}}\right\|,$ (7)
+where ${\bf X}^{t_{i},t_{i}}_{1}$} and ${\bf X}^{t_{i},t_{i}}_{2}$} indicate the output from Head1 and Head2, respectively. $\bar{{\bf X}}^{t_{i},t_{i}}$ is the pointmap groundtruth.
+3D Reconstruction.
+Assisted by the temporal motion module, redundant post-process operations such as global alignment can be omitted, allowing the reconstructed 3D point cloud to be obtained in a feed-forward manner.
+As shown in the bottom part of Fig.5, the keyframe is set to the last frame ${\bf I}^{t_{T}}$ within the temporal window of length $T$ and is fed to Head1 with a set output of $\{{\bf X}^{t_{T},t_{T}},{\bf X}^{t_{T},t_{T}},...,{\bf X}^{t_{T},t_{T}}\}$.
+All the reference frames are input to the Head2 so the target pointmaps $\{{\bf X}^{t_{1},t_{T}},{\bf X}^{t_{2},t_{T}},...,{\bf X}^{t_{T},t_{T}}\}$ are aligned under the coordinate system of the keyframe.
+The temporal consistency loss $\mathcal{L}_{\text{t}}$ is:
+$\mathcal{L}_{\text{t}}=\frac{1}{T}\sum_{i=1}^{T}\left\|\frac{{{\bf X}^{t_{T},t_{T}}}}{z^{T}_{1}}-\frac{{\bar{{\bf X}}^{t_{T},t_{T}}}}{\bar{z}_{1}^{T}}\right\|+\left\|\frac{{{\bf X}^{t_{i},t_{T}}}}{z_{2}^{T}}-\frac{{\bar{{\bf X}}^{t_{i},t_{T}}}}{\bar{z}_{2}^{T}}\right\|$ (8)
+We further freeze the parameters in Decoder1 and Decoder2 when training the temporal downstream tasks at the second stage.
+In our work, the temporal window length $T$ is set to 12.
+Additional explorations on the temporal length can be found in Sec.4.
 Sintel [4] BONN [30] KITTI [13]
-Alignment Method Optim. Onl. Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$ <$1.25\uparrow$
+Alignment Method Optim. Onl. Abs Rel $\downarrow$ $\delta$$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$ $1.25\uparrow$
 Per-sequence scale DUSt3R-GA [44] 0.656 45.2 0.155 83.3 0.144 81.3
@@
 Table 1: Video depth evaluation.
-We report scale-invariant depth and scale &\&& shift invariant depth accuracy on Sintel [4], Bonn [28], and KITTI [12] datasets.
+We report scale-invariant depth and scale &\& shift invariant depth accuracy on Sintel [4], Bonn [30], and KITTI [13] datasets.
 Methods requiring global alignment are marked “GA”, while “Optim.” and “Onl.” indicate optimization-based and online methods, respectively.
@@
 The specific number and the usage ratio of each dataset can be found in the supplementary materials.
-All of these datasets are synthetic and come with pixel-accurate ground truth depth, as well as camera intrinsics and extrinsics.
-They feature diverse dynamic scene types: indoor and outdoor.
-Among them, PointOdyssey and DynamicReplica have additional 2D trajectory annotations which can be used to construct pointmap matching ground truth following Eq.
+All datasets include pixel-accurate ground truth depth, as well as camera intrinsics and extrinsics, and encompass a wide variety of dynamic scenes across both indoor and outdoor environments.
+Among them, PointOdyssey and DynamicReplica have additional 2D trajectory annotations for dynamic objects which can be used to construct pointmap matching ground truth following Eq.
 2.
-Thus, all datasets contribute to geometry supervision for training pointmaps in Head1 and Head2 and we use only PointOdyssey, DynamicReplica, and TartanAir datasets to train our proposed pointmap matching head.
-The training of POMATO follows a two-stage process.
-In the first stage, we utilize pair-wise data to establish the fundamental capabilities of both geometry and matching.
-All parameters in the decoders and each DPT head are learnable.
-In the second stage, we introduce the temporal motion module and freeze parameters in the decoder.
-The same datasets are used in both stages.
+All datasets are used to supervise geometry learning on Head1 and Head2, while only PointOdyssey, DynamicReplica, and TartanAir are used to train the proposed pointmap matching head.
 Training and inference details.
 Our model architecture is based on the publicly available DUSt3R [52] model, utilizing the same backbone consisting of a ViT-Large encoder and a ViT-Base decoder.
-To maximize the benefits of MonST3R’s geometry estimation ability in dynamic scenes, we initialize the model weights using the publicly available MonST3R checkpoint.
-For the newly introduced pointmap matching head, we initialize the head weights from the Head2 weight of the MonST3R.
-We train our network for 10 epochs with a cosine schedule and the initial learning rate is set to 0.0001.
-The batch size for the first stage of pairwise image training is 16 on 4 A100 (40G) GPUs.
-When training with the temporal motion module, the batch size is set to 4 with a temporal window length of 12.
+To fully leverage MonST3R’s geometry estimation capabilities in dynamic scenes, we initialize our model using the publicly available MonST3R checkpoint.
+For the newly introduced pointmap matching head, we initialize its weights from the pretrained Head2 weights of MonST3R.
+The temporal motion module is initialized following [14].
+We train our network for 10 epochs with a cosine learning rate schedule, with an initial learning rate of 1e-4.
+In the first stage, which involves pairwise training, we use a batch size of 16 on 4 A100 GPUs (40G).
+In the second stage, where the temporal motion module is introduced, the batch size is set to 4 with a fixed temporal window length of 12.
+During each training iteration, we randomly sample a downstream task—3D point tracking, video depth estimation, or 3D reconstruction—to construct the input pairs and apply the corresponding loss function.
 4.2 Video Depth Estimation
@@
 Thus, we can measure the per-frame depth quality and inter-frame depth consistency.
-We employ our proposed motion module for video depth estimation as described in Sec.3.4 and compare our method against several variants of DUSt3R, including DUSt3R [42], MAST3R [24], MonST3R [50], Spann3R [39], and CUT3R [41].
-Some of these methods [42, 24, 50] rely on global alignment (GA), which assumes a static environment.
-While GA benefits multi-view consistency in static regions, it can potentially degrade the representation of dynamic elements and is computationally expensive.
-In contrast, POMATO can perform online video depth inference with the introduced temporal fusion module and achieve much faster inference speed.
+We employ our proposed motion module for video depth estimation in a feed-forward manner as described in Sec.3.5 and compare our method against several variants of DUSt3R, including DUSt3R [44], MAST3R [26], MonST3R [52], Spann3R [41], and CUT3R [43].
+Given 6 frames of 288×512 on an NVIDIA 4070 GPU, POMATO reconstructs the 3D point cloud in 0.7 seconds, whereas global alignment-based methods such as MonST3R require 5.8 seconds.
 As shown in Tab.
@@
 PointOdyssey [54] ADT [31] PStudio [20] Average
-Method T-12 T-24 T-12 T-24 T-12 T-24 T-12 T-24
-SpatialTracker [45] 21.24 23.33 22.34 22.87 32.94 32.81 25.51 26.34
-DUSt3R [42] 20.27 21.01 30.77 28.64 10.91 6.39 20.65 18.68
-MASt3R [24] 17.36 18.29 27.18 25.51 12.83 7.73 19.12 17.17
-MonST3R [50] 28.50 29.29 29.82 28.42 18.23 10.81 25.52 22.84
-POMATO 34.19 35.75 33.41 30.72 27.05 23.81 31.55 30.09
+Method L-12 L-24 L-12 L-24 L-12 L-24 L-12 L-24
+SpatialTracker [47] 20.46 20.71 21.64 20.67 30.41 25.87 24.17 22.42
+DUSt3R [44] 19.03 19.03 29.02 25.55 9.72 6.50 19.26 17.03
+MASt3R [26] 16.58 17.35 27.36 26.46 11.78 8.09 18.57 17.30
+MonST3R [52] 27.31 27.92 28.30 26.13 16.50 11.06 24.03 21.70
+POMATO 33.20 33.58 31.57 28.22 24.59 19.79 29.79 27.20
 Table 2: 3D tracking evaluation.
 We report the APD metric to evaluate 3D point tracking on the PointOdyssey [54], ADT [31], and PStudio [20] datasets.
-T-12 and T-24 indicate tracking within the temporal length of 12 frames and 24 frames, respectively.
-POMATO achieves remarkable performance even compared with the specialized model SpatialTracker [45].
-* indicates the camera intrinsic is required.
+L-12 and L-24 indicate tracking within the temporal length of 12 frames and 24 frames, respectively.
 Figure 6: Qualitative comparison of dynamic scenes. Compared to MonST3R, our POMATO can provide more reliable motion masks, 3D point tracking, and reconstruction performance.
@@
 The APD metric serves as a direct measure of the accuracy of the predicted tracking.
-We reformulate the datasets and project all the query points within a testing sequence to the first frame.
-We report tracking results on a length of both 12 and 24 frames.
+We reformulate the datasets and project all the query points within a temporal window to the first frame.
+We report tracking results on the length of 12 and 24 frames.
 As shown in Tab.2, our POMATO achieves the best performance on both PointOdyssey and ADT datasets.
-Additionally, our method demonstrates superior generalization, as reflected by the highest average APD metric.
 It’s worth mentioning that SpatialTracker [47] is a state-of-the-art network tailored for 3D point tracking with ground truth camera intrinsic as additional input data.
-POMATO surpasses it on two datasets and improves the average APD metric by 23.7% and 14.2% for 12 frames and 24 frames, respectively.
-In contrast, DuST3R-based methods struggle with ambiguous matching representations, resulting in imprecise tracking performance in dynamic scenarios.
-Video Depth Tracking
-Temporal Length Sintel [4] Bonn [28] KITTI [12] ADT [29] PStudio [18] PointOdyssey [52]
+POMATO surpasses it on two datasets and improves the average APD metric by 23.3% and 21.4% for 12 frames and 24 frames, respectively.
+For DUSt3R-based methods, we use the output of Head2 as tracking results.
+Obviously, the ambiguous matching representation limits its capability to handle this fine-grained 3D reconstruction task in dynamic scenes.
+Video Depth Tracking (12 Frames)
+Temporal Length Sintel [4] Bonn [30] KITTI [13] PointOdyssey [54] ADT [31] PStudio [20]
 […]
-Pair-wise 0.548 46.2 0.087 94.0 0.113 89.5 31.35 25.20 33.21
-6 frames 0.436 51.3 0.076 95.9 0.085 93.5 32.63 26.93 33.88
-12 frames 0.416 53.6 0.075 96.1 0.086 93.3 33.41 27.05 34.19
+Pair-wise 0.548 46.2 0.087 94.0 0.113 89.5 32.06 29.87 23.10
+6 frames 0.436 51.3 0.076 95.9 0.085 93.5 32.69 30.93 24.52
+12 frames 0.416 53.6 0.075 96.1 0.086 93.3 33.20 31.57 24.59
 Table 3: Ablation study on the temporal motion module.
-The introduction of the temporal motion module brings a significant improvement compared with the fundamental model trained on pairwise images.
-As the temporal window length enlarges from 6 frames to 12 frames, we obtain an overall consistent improvement for both video depth estimation and 3D point tracking tasks.
+The introduction of the temporal motion module brings a significant improvement.
+As the temporal window length enlarges from 6 frames to 12 frames, we obtain an overall consistent improvement.
 TUM [37] Bonn [30]
@@
 CUT3R [43] 0.023 0.016 0.510 0.028 0.033 2.569
-POMATO 0.020 0.010 0.509 0.037 0.016 1.837
-Table 4: Pose estimation.
-* indicates using an off-the-shelf optical flow model to get the motion mask and the pseudo 2D optical flow ground truth.
-Our method achieves an overall best performance and improves the RPE rot metric significantly.
-Figure 7: Effectiveness of our motion mask estimation in 3D reconstruction. Without explicitly filtering out the motion area, both pose and geometry estimation will be degenerated.
+POMATO 0.020 0.010 0.509 0.037 0.016 1.782
+Table 4: Pose estimation. Our method achieves an overall best performance and improves the RPE rot metric significantly.
+Figure 7: Effectiveness of our pointmap matching head. Without explicitly filtering out the motion area, both pose and geometry estimation will be degraded.
 4.4 Camera Pose Estimation
-We conduct pose estimation experiments on Bonn  [28] and TUM [35] datasets which include dynamic moving objects, and compare our method with the DUSt3R-related works along with the recently proposed CUT3R[41].
-Global alignment is utilized for the evaluation.
-The sampling stride is set to 5 on the Bonn dataset and 3 on the TUM dataset.
-We evaluate over 40 frames and report the results in Tab.
+Following DUSt3R-based methods, we perform global alignment with the model trained in the first stage on the Bonn [30] and TUM [37] datasets.
+The sampling stride is set to 5 for the Bonn dataset and 3 for the TUM dataset.
+Compared with optical-flow assisted global alignment in MonST3R, the dynamic mask is computed according to Eq.
+5 while the 2D pseudo label is replaced by projecting the pointmap matching results to 2D coordinates with estimated camera intrinsic.
+Absolute Translation Error (ATE), Relative Translation Error (RPE trans), and Relative Rotation Error (RPE rot) are reported.
+The evaluation results over 40 frames are reported in Tab.
 4.
-Three metrics are reported: Absolute Translation Error (ATE), Relative Translation Error (RPE trans), and Relative Rotation Error (RPE rot).
-Our method achieves the best overall performance.
-Notably, POMATO operates without relying on any auxiliary modules for dynamic object estimation, thereby avoiding potential domain gaps across modules and preventing error accumulation.
-In particular, POMATO significantly improves the RPE-rot metric, surpassing MonST3R by 55.4% and 13.3% on the TUM and Bonn datasets, respectively.
-We present a visualization to demonstrate the effectiveness of our dynamic mask estimation in 3D reconstruction, as mentioned in Sec.3.3.
-Without explicitly identifying dynamic regions, both camera pose and geometry estimation suffer from significant degeneration.
+Notably, POMATO obtains an overall state-of-the-art performance and significantly improves the RPE-rot metric, surpassing MonST3R by 55.4% and 13.3% on the TUM and Bonn datasets.
 4.5 Ablation Study
-Extensive ablation experiments are conducted on video depth estimation and 3D point tracking tasks to validate the effectiveness of learning temporal motions on video sequence input.
-We report three models trained with pairwise images, a shorter temporal window length of 6 frames, and the default temporal window length of 12 frames.
-As shown in Tab.
-3, the introduction of the temporal motion module introduces significant improvements across all datasets, emphasizing the importance of consistency among different frames.
-When enlarging the temporal window length from 6 frames to 12 frames, the video depth estimation obtains a further improvement on both Sintel and Bonn datasets.
-For 3D point tracking, the enlarged temporal window length brings a consistent enhancement across all three testing datasets.
+Bonn [30] PointOdyssey [54] ADT [31] PStudio [20]
+Method ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$ APD $\uparrow$ APD $\uparrow$ APD $\uparrow$
+W/O Head3 0.040 0.015 1.721 29.10 29.62 16.94
+W/ Head3 0.037 0.016 1.782 32.06 29.87 23.10
+Table 5: Ablation study on the effectiveness of the pointmap matching head. The comparisons are reported on the pose estimation and 3D point tracking tasks.
+We conduct extensive ablation studies to evaluate the effectiveness of the temporal motion module and the proposed pointmap matching head.
+As shown in Table 3, we report results for three models: one trained with only pairwise images (first-stage training), one using a shorter temporal window of 6 frames, and another using the default temporal window length of 12 frames.
+Incorporating temporal consistency yields substantial improvements across all datasets for video depth estimation and 3D point tracking.
+Further improvement is achieved when the temporal window length increases from 6 frames to 12 frames.
+In Table 5, we evaluate the effectiveness of the pointmap matching head.
+While it introduces only a modest improvement in the ATE metric, we attribute this to the limited motion and minimal viewpoint variation in the indoor evaluation dataset.
+As illustrated in Fig.
+7, under challenging in-the-wild conditions with significant motion and rapid viewpoint changes, removing the pointmap matching head introduces ambiguity in explicit rigid transformation estimation, resulting in a clear degradation in performance.
+To further demonstrate the impact of the pointmap matching head on 3D point tracking, we conduct tracking experiments over 12 frames using the pairwise input setup.
+Clearly, removing the pointmap matching head (using only Head2) leads to an inevitable performance drop, emphasizing explicit correspondence modeling for reliable long-term tracking.
 5 Discussion and Conclusion
-We introduce POMATO, a unified framework designed for geometry estimation and motion understanding in dynamic scenes.
-Leveraging our proposed pointmap matching head, our method can effectively distinguish moving regions, thereby mitigating the interference caused by dynamic objects.
-The introduced temporal motion module enhances the learning of temporal motions across different frames, improving the scale consistency and boosting the performance in tasks where geometry and matching are critical, most notably, 3D point tracking.
-Moving forward, we aim to explore methods to scale up our training with more matching data and further improve the 3D reconstruction performance.
+We introduce POMATO, a unified framework for geometry estimation and motion understanding in dynamic scenes.
+By leveraging the proposed pointmap matching head, our method effectively distinguishes moving regions, thereby mitigating the interference introduced by dynamic objects.
+The temporal motion module further facilitates the learning of temporal dynamics across frames, enhancing scale consistency and improving performance in tasks where both geometry and matching are critical—most notably, 3D point tracking.
+The downstream temporal tasks including 3D point tracking, video depth estimation, and 3D reconstruction can be easily applied in a feed-forward manner.
+In future work, we plan to scale up training with more dynamic reconstruction and matching datasets to further enhance 3D reconstruction and tracking performance.
+Acknowledgement. This work was supported by the National Natural Science Foundation of China (No. 62206244)
 Supplementary Material
@@
 In POMATO, for an image pair $\{{\bf I}^{i},{\bf I}^{j}\}$, the dynamic mask ${\bf D}^{j,i}$ is calculated by comparing the difference between ${\bf X}^{j,i}$ and ${\bf X}^{j,i}_{m}$:
-${\bf D}^{j,i}=||{\bf X}^{j,i}_{m}-{\bf X}^{j,i}||>\alpha,$ (7)
-where $\alpha$ is a dynamic threshold defined as $3\times\text{median}(||{\bf X}^{j,i}_{m}-{\bf X}^{j,i}||)$.
+[…]
+[…]
 Given the updated camera intrinsic $\tilde{K}$ after an iteration of optimization, the target matching 2D coordinates ${\bf F}^{j,i}_{m}\in{\mathbb{R}}^{H\times W\times 2}$ can be calculated as ${\bf F}^{j,i}_{m}=p(\tilde{\bf K}{\bf X}^{j,i}_{m})$ where $p$ is a mapping from 3D camera coordinates to 2D pixel coordinates.